# GestureGPT: Toward Zero-shot Free-form Hand Gesture Understanding with Large Language Model Agents

### v4

@@ Abstract.
Current gesture interfaces typically demand users to learn and perform gestures from a predefined set, which leads to a less natural experience. Interfaces supporting user-defined gestures eliminate the learning process, but users still need to demonstrate and associate the gesture to a specific system function themselves. We introduce GestureGPT, a free-form hand gesture understanding framework that does not require users to learn, demonstrate, or associate gestures. Our framework leverages the large language model’s (LLM) astute common sense and strong inference ability to understand a spontaneously performed gesture from its natural language descriptions, and automatically maps it to a function provided by the interface. More specifically, our triple-agent framework involves a Gesture Description Agent that automatically segments and formulates natural language descriptions of hand poses and movements based on hand landmark coordinates. The description is deciphered by a Gesture Inference Agent through self-reasoning and querying about the interaction context (e.g., interaction history, gaze data), which a Context Management Agent organizes and provides. Following iterative exchanges, the Gesture Inference Agent discerns user intent, grounding it to an interactive function. We validated our conceptual framework under two real-world scenarios: smart home controlling and online video streaming. The average zero-shot Top-5 grounding accuracies are 83.59% for smart home tasks and 73.44% for video streaming. We also provided an extensive discussion of our framework including model selection rationale, generated description quality, generalizability etc.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Gestural inputccs: Human-centered computing User interface management systems
Figure 1. A user aims to adjust the light’s brightness, assuming GestureGPT is successfully implemented, so he (a) triggers the system with a raised right hand above the chest, followed by looking at the light and performing a ‘pinch and move up’ gesture. The (b) input camera captures this action in real-time video streaming. (c) The Gesture Inference Agent then processes the gesture by (d) filtering key frames from the video and extracting hand landmarks for (e) generating the gesture state matrix. Subsequently, (f) the agent generates a description of the gesture’s pose and movement. Upon activated by the description, the (g) Gesture Inference Agent analyzes this description to understand which function the gesture maps to and (h) assesses the confidence level of the result. If the confidence is deemed insufficient, the agent requests context information from the (i) Context Management Agent, which retrieves relevant data from the (j) context library containing all available context in the current environment, to inform the inquiry. According to the inquiry, as context information such as ”Gaze: user is looking at the Light” and ”History: the user opened the light at 19:00” is incorporated, the Gesture Inference Agent (k) concludes that the function designated by the gesture is ‘Brightness Increase’ of the light. The function could then be triggered to finish the interaction process.
@@ 1. Introduction
Gestures express human intent in an intuitive and immediate manner, enables natural human computer interaction with low cognitive load (Pavlovic et al., 1997; Xia et al., 2022). Most existing gesture interfaces only support pre-defined gestures. Within the pre-defined gesture set, such interfaces can achieve a very high gesture classification accuracy (Wah Ng and Ranganath, 2002; Berman and Stern, 2012). However, it takes great efforts to learn and memorize different gestures and their mappings to different system functions (Gao et al., 2023a; Ejaz et al., 2019), which is especially true with a large gesture set. The mapping between the gesture and the function is also fixed, which cannot be easily expanded and may contradict the user’s preference  (Delamare et al., 2019; Vatavu, 2012; Nacenta et al., 2013).
@@
To address such limitations of pre-defined gestures, researchers propose gestural interfaces that support user-defined gestures (Lü and Li, 2012; Xu et al., 2022; Wang et al., 2021). Users can define their own gestures for each function, and they only need to demonstrate the gesture several times, which eliminates the gesture learning effort and offers greater system flexibility. However, each user needs to define and demonstrate her own gesture, as well as associate it with a system function (Mo et al., 2021; Speicher and Nebeling, 2018). Users also still need to memorize their own gestures for different functions, which may lead to frustration using the interface. Such constraints deteriorate the natural interaction experience and hinder the wide adoption of gestural interfaces, which leads to Norman’s famous argument that ”Natural User Interfaces are Not Natural” (Norman, 2010). While significant advances have been made in gesture recognition technologies over the years, this naturalness challenge of gestural interfaces persists.
@@
A natural gesture interface requires end-to-end gesture understanding rather than merely gesture recognition. Users do not need to learn, memorize, or demonstrate gestures. They can just perform gestures as they see fit based on their common sense, which involves the semantics of the function and their daily interaction experience with humans and machines. The association between the gesture and the function is also automatically executed by the interface by considering both the gesture and the interaction context. However, gesture understanding is inherently more challenging than gesture recognition.
@@
Towards a natural gesture understanding interface, we propose GestureGPT, an LLM-based hand gesture interface framework that understands natural free-form hand gestures and automatically associates the gesture with the intended function. The core idea of GestureGPT is to leverage LLM’s rich common sense to recognize gestures and understand the interaction context, as well as its strong inference capability to map the gesture to the intended function. However, it still involves plenty of efforts to build such an LLM-based gesture understanding framework. For example, gestures and context information need to be transformed and formatted for LLMs to understand. There also need to be a mechanism for robust and thorough synthesizing of all relevant information without forgetting any of it.
@@
To that end, we design a triple-agent framework to generate gesture description, manage the context, and infer the intended function, which involves a Gesture Description Agent, a Gesture Inference Agent and a Context Management Agent, all based on LLMs. 1) Gesture Description Agent describes hand gestures in natural languages to facilitate the powerful text-based LLMs reading and understanding of gestures. A set of specially designed and tuned rules transform hand poses and movements into discreet states (e.g., the bending state of each finger) based on visually-extracted hand landmarks, triggered by raising hand to the chest level. The agent then synthesizes a general description of the gesture based on a matrix formed with the states. 2) Gesture Inference Agent analyzes the description and initiates conversation with Context Management Agent to inquire about the context information. After multi-round dialogues, it infers which interactive function the user intents to activate. 3) Context Management Agent answers questions of Gesture Inference Agent by referring to a context library, which includes different types of context information like gaze point, interaction history, etc.  The context library is organized in the form of JSON files.
@@
GestureGPT offers a number of advantages that address the challenges and limitations of previous gestural interfaces. 1) GestureGPT allow users to perform natural free-form hand gestures, which do not need to resemble those in a pre-defined gesture set or those previously defined by users. This eliminates any learning, memorizing, or demonstration effort. 2) GestureGPT automatically associates gestures with corresponding interface functions by step-by-step inference based on both the gesture description and the interaction context. Such a novel LLM-based structure successfully addresses gesture’s ephemeral nature and close relationship with context (Norman, 2010). Additionally, the integration of context further enhances the accuracy of gesture mapping. 3) GestureGPT analyzes spatial coordinates of hand landmarks to generate gesture descriptions, making it independent of view angles and even modalities. For example, hand landmarks can also be reconstructed by wearable sensors (Jiang et al., 2020; Zhang et al., 2018b). Such flexibility makes it easy to adapt GestureGPT for a wide range of applications. The natural language descriptions also preserve user privacy and reduce data transmission load, which are important for interaction interfaces.
@@
Describing nuanced states and movements of fingers are key to accurately understand hand gestures. Even though it is a difficult task for current large vision models, our hand landmarks based method captures finger states accurately and thoroughly by first calculating a set of discreet finger states based on rules, then use an LLM to synthesize a gesture summary. The parameters for the rules are tuned on third-view public gesture image datasets, and test on both third-view dataset (HaGRID (Kapitanov et al., 2024), 38576 test samples) and first-view dataset (EgoGesture (Zhang et al., 2018a), around 5000 test samples). The overall error rate is 2.3% on the third-view dataset and 6.3% on the first-view dataset. The synthesized summary is rated by two gesture experts and achieved a 3.51($\text{std}=1.14$) on a 5-point Likert scale (1 being almost none and 5 being almost all of the description is correct and relevant to the gesture).
@@
To evaluate our entire triple-agent framework, we ran experiments under two realistic interaction scenarios to test boundaries of our framework: smart home IoT device control with AR headset (first-person view with 18 functions) and online video streaming on a desktop PC (third-person view with 66 functions). The highest zero-shot Top-1/Top-5 gesture grounding accuracy is 44.79%/83.59% for smart home control and 37.50%/73.44% for video streaming. We report accuracy at different levels to better understand and demonstrate the framework’s performance and boundaries. The results shows great potential for our framework, which is the first zero-shot free-form gesture understanding framework that does not require any learning, memorization, demonstration, and association efforts as to our best knowledge.
GestureGPT does not yet offer a practical interface that can be used today-primarily due to the slow inference speeds of LLM systems, which in our evaluation average 227 seconds per task. As a new gesture interface paradigm, however, GestureGPT framework serves as a foundation for further innovations. For example, future advances on edge-side large language models or large multi-modal models fine-tuned for gesture understanding tasks can greatly reduce system response delay, approaching a practical interface in the end.
@@
    We proposed and evaluated the first framework for automatic free-form hand gesture understanding, as to our best knowledge. Our framework leverages LLMs to mimic human gesture understanding process.
@@
    We designed a set of gesture description rules based on hand landmarks to thoroughly and accurately capture states and movements of fingers, which has similar performance with SOTA large multi-modal modal GPT-4o.
@@
    We carefully crafted prompts for each agent and evaluated the effects of different contexts on our framework’s performance. Such insights are invaluable for future context-aware agent-based gesture understanding work.
@@ 2.2. Natural Free-form Gesture Understanding
Gestural interfaces working with pre-defined gesture set demand large annotated dataset, and the confinement to pre-defined gestures also hampers the naturalness of interaction (Wexelblat, 1995). The conflicts in different system designs further exacerbate user adaptation challenges across platforms. User-defined gestural interfaces mitigates the learning burden by allowing users to define their own gestures. Only several demonstrations of the gesture are necessary for the system to learn new gestures with the help of advanced few-shot learning algorithms (Xu et al., 2022). Gesture Coder (Lü and Li, 2012) allows the user to demonstrate a gesture, and, instead of defining a gesture name for it, the user directly associate it to a designated function with auto-generated recognizer. But in these works, users still need to demonstrate and memorize the gesture while ensuring its distinctiveness from existing gesture commands. Recent advances in zero-shot learning make it possible to recognize gesture classes unseen during training, but it requires manually introducing prior knowledge (Madapana and Wachs, 2019, 2018) or representative samples (Wu et al., 2021) of unseen class.
@@
Nevertheless, all aforementioned approaches are still restricted by a finite set of distinguishable gestures defined either by the interface designers or users. This results in a mismatch between the fixed gesture set and the flexible gestures human performs understand different scenarios, which greatly restricts the inherent expressiveness of gesture. Thus, both pre-defined and user-defined gestural interfaces actually require users to adapt to it, rather than the reverse (Wexelblat, 1995). Free-from gestural interfaces, on the other hand, do not have such restrictions. For example, Gesture Avatar (Lü and Li, 2011) allows any form of drawing gestures on a screen. The input gesture is associated with a UI element based on their appearance resemblance, providing intuitive interaction experience. However, there are still gaps in research of free-form hand gesture understanding. Compared with screen gestures, hand gestures are more complicated in form, do not necessarily resemble UI elements in appearance, and can change under different scenarios even for the same function. To tackle such challenges, our framework relies on LLMs to understand the semantic meanings of both the hand gestures and the system functions to ensure correct gesture-function mapping. An overview of current gesture-based interaction systems and the placement of our work is shown in Figure 2.
@@ 3. Method
GestureGPT system has a triple-agent framework to efficiently manage gesture description, context, and inference tasks, thereby enhancing system flexibility and scalability while addressing the shortcomings of single or dual agent systems, consists of: (1) Gesture Description Agent, generating descriptions of gestures from videos; (2) Context Management Agent, handling interaction context; and (3) Gesture Inference Agent, synthesize information to infer gesture intentions through dialogue and reasoning.
The workflow initiates with the transformation of user gestures, captured by an RGB camera, into natural language description. The Gesture Description Agent first uses a set of rules to transform the gesture into a ”Gesture State Matrix”, which delineates hand and finger states over time. The agent then relies on this matrix to produce a summary of gestures, which is forwarded to the Gesture Inference Agent. Upon receiving the gesture description, the Gesture Inference Agent engages in multi-round dialogue with the Context Management Agent. This involves identifying and soliciting relevant context information from a context library managed by the Context Management Agent. Through iterative dialogue, the Gesture Inference Agent gains a comprehensive understanding of the interactive scenario and completes a dynamic mapping between the gesture and possible functions.

### v5

@@ Abstract.
Existing gesture interfaces only work with a fixed set of gestures defined either by interface designers or by users themselves, which introduces learning or demonstration efforts that diminish their naturalness. Humans, on the other hand, understand free-form gestures by synthesizing the gesture, context, experience, and common sense. In this way, the user does not need to learn, demonstrate, or associate gestures. We introduce GestureGPT, a free-form hand gesture understanding framework that mimics human gesture understanding procedures to enable a natural free-form gestural interface. Our framework leverages multiple Large Language Model agents to manage and synthesize gesture and context information, then infers the interaction intent by associating the gesture with an interface function. More specifically, our triple-agent framework includes a Gesture Description Agent that automatically segments and formulates natural language descriptions of hand poses and movements based on hand landmark coordinates. The description is deciphered by a Gesture Inference Agent through self-reasoning and querying about the interaction context (e.g., interaction history, gaze data), which is managed by a Context Management Agent. Following iterative exchanges, the Gesture Inference Agent discerns the user’s intent by grounding it to an interactive function. We validated our framework offline under two real-world scenarios: smart home control and online video streaming. The average zero-shot Top-1/Top-5 grounding accuracies are 44.79%/83.59% for smart home tasks and 37.50%/73.44% for video streaming tasks. We also provide an extensive discussion that includes rationale for model selection, generalizability, and future research directions for a practical system etc.
@@
copyright: rightsretainedjournalyear: 2024doi: 10.1145/3698145journal: PACMHCIjournalvolume: 8journalnumber: ISSarticle: 545publicationmonth: 12ccs: Human-centered computing Gestural inputccs: Human-centered computing User interface management systems
Figure 1. This scenario illustrates a system implemented based on GestureGPT.
@@ 1. Introduction
Gestures express human intent intuitively and promptly, enabling natural human-computer interaction with low cognitive load (Pavlovic et al., 1997; Xia et al., 2022). Most existing gesture interfaces support only predefined gestures, which can achieve a very high gesture classification accuracy (Wah Ng and Ranganath, 2002; Berman and Stern, 2012). However, it requires significant effort to learn and memorize different gestures and their respective mappings to system functions, particularly with a large set of gestures (Gao et al., 2023a; Ejaz et al., 2019). The gesture-function mapping is also fixed, which cannot be easily expanded and may contradict users’ preferences (Delamare et al., 2019; Vatavu, 2012; Nacenta et al., 2013).
@@
To address the limitations of predefined gestures, researchers have proposed gestural interfaces that support user-defined gestures (Lü and Li, 2012; Xu et al., 2022; Wang et al., 2021). Users can define their own gestures for each function with only a few demonstrations, thus eliminating the learning effort and enhancing system flexibility. However, each user must design and demonstrate their own gesture and associate it with a system function (Mo et al., 2021; Speicher and Nebeling, 2018). Users also still need to memorize their own gestures, which may lead to frustration using the interface, especially when there are a large number of functions. Such constraints degrade the natural interaction experience and hinder the broad adoption of gestural interfaces, leading to Norman’s famous argument that “Natural User Interfaces are Not Natural” (Norman, 2010). While significant advances have been made in gesture recognition technologies over the years, this naturalness challenge of gestural interfaces persists. To address the pursuit of more natural gestural expressions, users expect high-frequency gestures to seamlessly link various functions (e.g., common touchscreen gestures like zooming in/out, or typical semantic gestures like thumbs-up and rock-and-roll) without being confined to a limited set of expressions or frequently needing to customize their own (Jian-peng et al., 2019; Li et al., 2012). Additionally, as gestural interfaces become more widespread, users will likely experience significant improvements in both their familiarity with these expressions and the speed of their interactions (Stiegemeier et al., 2022; Murray and Häubl, 2010).
@@
To that end, future gestural interfaces require direct, end-to-end gesture understanding rather than merely gesture recognition. Users should not have to learn, memorize, or demonstrate specific gestures. Instead, they can perform gestures naturally, according to their understanding of the function’s semantics and their everyday interaction experiences with humans and machines. The interface automatically links the gesture to its corresponding function, considering both the gesture and the interaction context. This gesture understanding task is inherently more challenging than simple gesture recognition.
@@
We propose GestureGPT, an LLM-based paradigm that comprehends natural free-form hand gestures and automatically links them to their intended functions. The core idea of GestureGPT is to utilize the LLM’s rich common sense for recognizing gestures and understanding the interaction context, as well as its potent inference capabilities to map gestures to their intended functions. Constructing such an LLM-based gesture understanding framework demands considerable effort. For instance, gestures and context information must be transformed and formatted in a way that LLMs can process. Additionally, there needs to be a mechanism for robust and thorough synthesis of all relevant information without overlooking any details.
@@
To that end, we design a triple-agent paradigm that generates gesture descriptions, manages the context, and infers the intended function. This involves a Gesture Description Agent, a Gesture Inference Agent, and a Context Management Agent, all based on LLMs. 1) The Gesture Description Agent describes hand gestures in natural languages to facilitate the powerful text-based LLMs’ reading and understanding of gestures. A set of specially designed and tuned rules transforms hand poses and movements into discrete states (e.g., the bending state of each finger) based on visually extracted hand landmarks, triggered by raising the hand to the chest level. The agent then synthesizes a general description of the gesture based on a matrix formed with the states. 2) The Gesture Inference Agent analyzes the description and initiates a conversation with the Context Management Agent to inquire about the context information. After multiple rounds of dialogue, it infers which interactive function the user intends to activate. 3) The Context Management Agent answers questions from the Gesture Inference Agent by referring to a context library, which includes various types of context information, such as gaze points and interaction history, etc.  The context library is organized in JSON file format.
Figure 1 envisions a process in which a system, implemented through the GestureGPT concept, understands a user’s gesture. In this scenario, he (a) triggers the system with a raised right hand above the chest, followed by looking at the light and performing a “pinch and move up” gesture. The (b) input camera captures this action through real-time video streaming. (c) The Gesture Description Agent processes the gesture by (d) filtering key frames from the video and extracting hand landmarks to (e) generate the gesture state matrix. Subsequently, (f) Gesture Description Agent generates a description of the gesture’s pose and movement. Upon activation by the description, the (g) Gesture Inference Agent analyzes this description to determine which function the gesture maps to and (h) assesses the confidence level of the result. If the confidence is deemed insufficient, Gesture Inference Agent requests context information from the (i) Context Management Agent, which retrieves relevant data from the (j) context library containing all available context information in the current environment, to inform the inquiry. According to the inquiry, as context information such as “Gaze: user is looking at the light” and “History: the user turned on the light at 19:00” is incorporated, the Gesture Inference Agent (k) concludes that the function designated by the gesture is to “increase the brightness” of the light. The function could then be triggered to finish the interaction process.
@@
GestureGPT offers a number of advantages that address the challenges and limitations of previous gestural interfaces. 1) GestureGPT allows users to perform natural free-form hand gestures, which do not need to resemble those in a predefined gesture set or those previously defined by users. This eliminates the need for learning, memorizing, or demonstrating specific gestures. 2) GestureGPT automatically associates gestures with their corresponding interface functions through a step-by-step inference process based on both the gesture description and the interaction context. Such a novel LLM-based structure successfully addresses the ephemeral nature of gestures and their close relationship with context (Norman, 2010). Additionally, the integration of context further enhances the accuracy of gesture mapping. 3) GestureGPT analyzes spatial coordinates of hand landmarks to generate gesture descriptions, making it independent of view angles and even modalities. For example, hand landmarks can also be reconstructed by wearable sensors (Jiang et al., 2020; Zhang et al., 2018b). Such flexibility makes it easy to adapt GestureGPT for a wide range of applications. The natural language descriptions also preserve user privacy and reduce data transmission load, which are important for interaction interfaces.
@@
Describing nuanced finger states and movements is crucial for accurately understanding hand gestures. Our hand landmarks-based method captures finger states accurately and thoroughly by first calculating a set of discrete finger states based on rules, then using an LLM to synthesize a gesture summary. The rules’ parameters are tuned using third-person view public gesture image datasets and tested on both third-person view (HaGRID (Kapitanov et al., 2024), 38576 test samples) and first-person view datasets (EgoGesture (Zhang et al., 2018a), approximately 5000 test samples). The error rates are 2.3% for third-person views and 6.3% for first-person views, respectively. Two gesture experts rated the synthesized summary, which achieved a score of 3.51 ($\text{std}=1.14$) on a 5-point Likert scale, where 1 indicates that almost none of the description is correct and relevant, and 5 indicates that almost all of it is.
@@
To evaluate our framework, we conducted experiments under two realistic interaction scenarios: smart home IoT device control using an AR headset (first-person view with 18 functions) and online video streaming on a desktop PC (third-person view with 66 functions). The highest zero-shot Top-1/Top-5 gesture grounding accuracies achieved were 44.79%/83.59% for smart home control and 37.50%/73.44% for video streaming. We report accuracies at various levels to more comprehensively understand and demonstrate the framework’s performance and boundaries. The results demonstrate the significant potential of GestureGPT, which, to the best of our knowledge, is the first zero-shot free-form gesture understanding framework that requires no learning, memorization, demonstration, or association efforts.
However, GestureGPT is not yet a practical interface ready for everyday use, primarily due to the slow inference speeds of current LLM systems. Our current system averages 227 seconds per task in our evaluations, a delay caused by long prompts, multi-turn dialogues, and rate limits enforced by LLM cloud services. Instead, it introduces a new paradigm for gestural interfaces. To the best of our knowledge, GestureGPT is the first feasible framework for the inherently difficult task of understanding free-form gestures. Previous efforts in gesture recognition have not successfully addressed free-form gestures, let alone the more complex challenge of gesture understanding. GestureGPT, therefore, not only pioneers a new approach but also lays the foundational framework for future advancements. This framework can also be easily expanded to recognize a broader range of user intents through different modalities, such as facial expressions and body postures, among others. Technologies such as edge-side LLMs or specifically fine-tuned large multi-modal models could significantly reduce response times, paving the way for a more practical implementation of this paradigm in the future.
@@
    We proposed and evaluated the first framework that mimics human reasoning processes to achieve automatic free-form hand gesture understanding, as to our best knowledge.
@@
    We designed a set of gesture description rules based on hand landmarks to thoroughly and accurately capture states and movements of the hand, which has comparable performance to SOTA large multi-modal modal GPT-4o.
@@
    We carefully crafted prompts for each agent, analyzed the offline evaluation results, and generated various insights. Such insights are invaluable for future context-aware agent-based gesture understanding work.
@@ 2.2. Natural Free-form Gesture Understanding
Gestural interfaces working with pre-defined gesture set demand large annotated dataset, and the confinement to pre-defined gestures also hampers the naturalness of interaction (Wexelblat, 1995). The conflicts in different system designs further exacerbate user adaptation challenges across platforms. User-defined gestural interfaces mitigates the learning burden by allowing users to define their own gestures. Only several demonstrations of the gesture are necessary for the system to learn new gestures with the help of advanced few-shot learning algorithms (Xu et al., 2022). Gesture Coder (Lü and Li, 2012) allows the user to demonstrate a gesture, and, instead of defining a gesture name for it, the user directly associate it to a designated function with auto-generated recognizer. But in these works, users still need to demonstrate and memorize the gesture while ensuring its distinctiveness from existing gesture commands.
@@
Nevertheless, all aforementioned approaches are still limited by a finite set of distinguishable gestures defined by interface designers or users. This leads to a mismatch between the fixed gesture set and the flexible gestures humans perform in different scenarios, greatly restricting gesture expressiveness. Thus, both pre-defined and user-defined gestural interfaces require users to adapt to them, rather than the reverse (Wexelblat, 1995). Free-form gestural interfaces, however, do not have these limitations. For instance, Gesture Avatar (Lü and Li, 2011) enables any drawing gesture on a screen. The input gesture is linked to a UI element based on appearance resemblance, providing an intuitive interaction experience. Yet, gaps remain in the research of free-form hand gesture understanding. Compared with screen gestures, hand gestures are more complex, do not necessarily resemble UI elements, and can vary under different scenarios for the same function.
Recent advances in zero-shot learning make it possible to recognize unseen gesture classes (Madapana and Wachs, 2019, 2018; Wu et al., 2021). Madapana (Madapana and Wachs, 2019, 2018) proposed frameworks that define a set of attributes for gestures, recognizing unseen gestures by manually labeling these attributes for each unseen gesture. Similarly, Wu et al. (2021) introduced a prototype-based approach that requires the definition of “semantic prototypes” for unseen gesture categories. These methods explore to some extent the recognition of “unseen” gestures, enabling the system to recognize new gestures without additional training.
However, above methods focus solely on the task of gesture recognition. Even for the recognition task, they only identify key features of new gestures, not their semantic names. Gesture understanding, however, requires not only recognizing gestures but also mapping them to functions, which demands a comprehensive integration of complex contextual information (such as available functions, interaction history, and physical environment). Such analysis and reasoning necessitate mimicking human thought processes, a task that traditional machine learning models struggle to achieve.
To tackle such challenges, our framework relies on LLMs to understand the semantic meanings of both the hand gestures and the system functions to ensure correct gesture-function mapping, because LLMs are currently the most promising tools capable of human-alike inference and comprehensive context understanding. An overview of current gesture-based interaction systems and the placement of our work is shown in Figure 2.
@@ 3. Method
GestureGPT has a triple-agent framework to efficiently manage gesture description, context, and inference tasks, enhancing system flexibility and scalability while addressing the shortcomings of single or dual agent systems, consists of: (1) Gesture Description Agent, generating descriptions of gestures from videos; (2) Context Management Agent, handling interaction context; and (3) Gesture Inference Agent, synthesize information to infer gesture intentions through dialogue and reasoning.
The workflow initiates with the transformation of user gestures, captured by an RGB camera, into natural language description. The Gesture Description Agent first uses a set of rules to transform the gesture into a “Gesture State Matrix”, which delineates hand and finger states over time. The agent then relies on this matrix to produce a summary of gestures, which is forwarded to the Gesture Inference Agent. Upon receiving the gesture description, the Gesture Inference Agent engages in multi-round dialogue with the Context Management Agent. This involves identifying and soliciting relevant context information from a context library managed by the Context Management Agent. Through iterative dialogue, the Gesture Inference Agent gains a comprehensive understanding of the interactive scenario and completes a dynamic mapping between the gesture and possible functions.